You are called by Hope van Dyne to determine whether a new Pym technology for miniaturized food delivery is economically desirable. This project will require $50,000,000 in equipment to infuse food with modified Pym particles. Installing and calibrating the equipment to begin operations will cost an additional $4,500,000. This equipment is categorized as seven-years MACRS property. The project will bring a revenue of $35,000,000 in the first year. After that, the revenue will grow 5% every year (as more people take a liking to miniature food).

The main cost of this project (other than initial installation and calibration) is the salary of lab technicians, which in the first year adds up to $5,000,000. The salary of these workers will increase with the general inflation rate of 2% every year. In addition to this, rent and utilities cost an additional $5,000,000 every year. This cost is expected to stay constant. The project is expected to last ten years, at the end of which all the equipment can be sold for 10% of its original value.

Assuming a tax rate of 21% and an interest rate of 12%, compute income, cash flows, and net present worth over the project life.
Answer #1
A B C D K E F=A+C+D+E+K G=F*21% H=F-G-E I J=H*I Cost of Asset        5,45,00,000
Year Annual Revenue Investment in Equipment Salvage Value Salary Rent Depreciation Taxable Income Tax @21% After tax CF MARR @10% PV Year Rate Depreciation
Year 0                             -          -5,45,00,000                           -                             -                             -                         -                         -        -5,45,00,000     1.0000      -5,45,00,000 Year 1 14.29%           77,88,050
Year 1            3,50,00,000                           -                             -              -50,00,000            -50,00,000       -77,88,050      1,72,11,950         36,14,510          58,09,391     0.8929          51,86,956 Year 2 24.49%        1,33,47,050
Year 2            3,67,50,000                           -                             -              -51,00,000            -50,00,000    -1,33,47,050      1,33,02,950         27,93,620         -28,37,720     0.7972         -22,62,213 Year 3 17.49%           95,32,050
Year 3            3,85,87,500                           -                             -              -52,02,000            -50,00,000       -95,32,050      1,88,53,450         39,59,225          53,62,176     0.7118          38,16,691 Year 4 12.49%           68,07,050
Year 4            4,05,16,875                           -                             -              -53,06,040            -50,00,000       -68,07,050      2,34,03,785         49,14,795       1,16,81,940     0.6355          74,24,084 Year 5 8.93%           48,66,850
Year 5            4,25,42,719                           -                             -              -54,12,161            -50,00,000       -48,66,850      2,72,63,708         57,25,379       1,66,71,479     0.5674          94,59,845 Year 6 8.92%           48,61,400
Year 6            4,46,69,855                           -                             -              -55,20,404            -50,00,000       -48,61,400      2,92,88,051         61,50,491       1,82,76,160     0.5066          92,59,271 Year 7 8.93%           48,66,850
Year 7            4,69,03,347                           -                             -              -56,30,812            -50,00,000       -48,66,850      3,14,05,685         65,95,194       1,99,43,641     0.4523          90,21,491 Year 8 4.46%           24,30,700
Year 8            4,92,48,515                           -                             -              -57,43,428            -50,00,000       -24,30,700      3,60,74,386         75,75,621       2,60,68,065     0.4039       1,05,28,454
Year 9            5,17,10,941                           -                             -              -58,58,297            -50,00,000                       -        4,08,52,644         85,79,055       3,22,73,588     0.3606       1,16,38,180
Year 10            5,42,96,488                           -               50,00,000            -59,75,463            -50,00,000                       -        4,83,21,025     1,01,47,415       3,81,73,610     0.3220       1,22,90,881
Increased by 5% annually 10% of cost of equipment Increased by 2% NPV      2,18,63,640

Total Income is in Column F.

Cashflows are in Column H.

NPV = 21,863,640

Please Upvote and Support!!